Attachment_J._6.pdf

PDF 274 KB Posted

Attached to
Operations and Maintenance of the Robert E. Kamosa Transmitting Station (REKTS) Facility
Solicitation number
BBG50-R-15-0006
Issued by

About this file

J.6 Failure Logs January 21 2015 pg. 1-10.

View the file

Other files

File Type Posted
Questions_and_Answers_Final_for_BBG50-R-15-0006.docx DOCX document
SF-33.pdf PDF
NMI_REKTS_Drawings.zip ZIP file
Attachment_J._1.pdf PDF
BBG50-R-15-0006_FBO.pdf PDF
Attachment_J.5.pdf PDF
Attachment_J.4.pdf PDF
Attachment_J._7.pdf PDF
Attachment_J.9.pdf PDF
Attachment_J.3.pdf PDF
Attachment_J.2.pdf PDF
Show all 11

On GovTribe

Work with this file

  • Download the original file
  • Contacts named in this file
  • Similar government files
  • Ask GovTribe AI about this file

Text version

BROADCASTING BOARD OF GOVERNORS

INTERNATIONAL BROADCASTING BUREAU

Robert E. Kamosa Transmitting Station - SAIPAN

DAILY OPERATIONS LOG

Wednesday, January 21, 2015

BEGINNING OF RADIO DAY STATUS

ACTIVE DATABASE B14 Saipan Sched B14-009.MDB

TRANSMITTERS All transmitters operational in PLC automation mode

ANTENNA & SLEW All antenna and slew switches operational

REMARKS

PENDING WORK ORDERS

Time Time

Start End Description of Events

00:00 Beginning of Broadcast Day

On-Duty: Mgr. EWilliams, GSolis, JDavid, TVirrey, CGatharngeg, JDiaz

00:00 SAI-01 OFF-Air No Broadcast Schedule.

00:00 01:00 SAI-02 On Air (PLC Auto Mode) @ 15690 kHz, 270M/270°, RFA3-LAO.

00:00 SAI-03 OFF-Air No Broadcast Schedule.

00:01 Removed Antenna 340M interlock key from Unit 21 and attached grounding stick thru the feedline.

00:10 01:00 Conducted SAI-01 weekly Modulator Unit preventive maintenance and re-aligned ground microswitch lever.

01:00 Removed Antenna 270M interlock key from Unit 21 and attached grounding stick thru the feedline.

02:40 Removed Antennas 195M, 270M and 340M grounding sticks from the feedlines and returned interlock keys to Unit 21.

03:00 07:00 SAI-02 On-Air (PLC Auto Mode) @ 15665 kHz, 270M/300°, RFA1-MAND.

03:00 07:00 SAI-03 On-Air (PLC Auto Mode) @ 17690 kHz, 340M/325°, RFA1-MAND.

05:12 CSpencer arrived on site.

05:20 ETadios arrived on site.

05:43 MRahman arrived on site.

06:00 Evening Shift Starts – Beginning of Shift Status

On Duty: CSpencer & ETadios

Security Personnel: MRahman

06:05 GSolis departed site.

06:20 JDavid departed site.

07:05 Informed APingul at Tinang MCR that we will assume OpCon for maintenance.

07:20 Conducted weekly maintenance on SAI-02 Modulator Unit.

09:15 09:33 Tested transmitters per upcoming broadcast schedule.

Start End Decription of Event

Page 2 of 10 Attachment J. 7.docx

09:35 Transferred OpCon to JManalad at Tinang MCR.

11:00 12:00 SAI-01 On Air (PLC Auto Mode) @ 13685 kHz, 270M/270°, RFA3-LAO.

11:00 12:00 SAI-03 On Air (PLC Auto Mode) @ 9825 kHz, 340M/325°, R-CHIN.

12:00 14:00 SAI-03 On Air (PLC Auto Mode) @ 9825 kHz, 340M/310°, R-CHIN.

12:30 13:30 SAI-01 On Air (PLC Auto Mode) @ 11750 kHz, 270M/270°, RFA3-KHME.

12:30 13:30 SAI-02 On Air (PLC Auto Mode) @ 12105 kHz, 340M/310°, RFA2-BURM.

13:30 14:00 SAI-01 On Air (PLC Auto Mode) @ 11795 kHz, 270M/270°, RFA3-BURM.

13:30 14:30 SAI-02 On Air (PLC Auto Mode) @ 12105 kHz, 340M/310°, RFA3-BURM.

13:34 ABasher arrived on site.

14:00 Midnight Shift Starts – Beginning of Shift Status

On Duty: Unmanned

Security Personnel: ABasher

14:00 15:00 SAI-01 On Air (PLC Auto Mode) @ 11890 kHz, 270M/270°, RFA2-VIET.

14:00 15:00 SAI-03 On Air (PLC Auto Mode) @ 9910 kHz, 340M/310°, RL-1 RU.

14:05 Informed JManalad at Tinang MCR that midnight shift is unmanned.

Forwarded phone to Security Guard House.

14:03 MRahman departed site.

14:07 ETadios departed site.

14:07 CSpencer departed site.

21:19 ELabios arrived on site.

21:37 JDavid arrived on site.

21:22 Inquired from Nelson of Tinang MCR regarding midnight shift operations. He reported that 9355 khz didn’t go on air transfer the broadcast in SAI-02, with 4 minutes loss time.

22:00 Day Shift Starts – Beginning of Shift Status

On-Duty: Mgr. EWilliams, AGruspe, ELabios, JDavid, TVirrey, CGatharngeg, JDiaz

22:00 23:00 SAI-02 On Air (PLC Auto Mode) @ 9720 kHz, 270M/285°, RFA1-CANT.

Swapped from SAI-01 due to IPA power supply trouble-shooting.

22:01 ABasher departed site.

23:00 24:00 SAI-03 On Air (PLC Auto Mode) @ 9585 kHz, 340M/310°, RFA1-MAND.

24:00 End of Broadcast Day

TRANSMITTER POWER, FILAMENT & PLATE METER READINGS

Transmitter Power Filament Plate CUC Power

SAI-01 3506537.3 22:44 6:02 2801

SAI-02 01059863 24:00 13:58 CUC Water

SAI-03 01585801 24:00 16:09 262623.5

Page 3 of 10 Attachment J. 7.docx

OFF/ONS ENCOUNTERED

Transmitter Day Shift Swing Shift Mid Shift Daily Total

SAI-01 0 0 1 1

SAI-02 0 0 0 0

SAI-03 0 0 0 0

TOTAL 0 0 1 1

SCHEDULED HOURS, OUTAGES & RELIABILITY

Transmitter Schedule Hrs Off/On’s Outage Minutes % Reliability

SAI-01 11.50 1 0.04 99.42

SAI-02 8.00 0 0.00 100.00

SAI-03 16.00 0 0.00 100.00

TOTAL 35.50 1 0.04 99.81

AUTOMATION INTERVENTION LOG

Enter OFF-ON’s, Substitutions and Loss Times Here

17:05 – 17:09 Off air due to PLC Error indication according to Tinang MCR.

17:09 – 22:00 SAI-01, 9355 kHz, 270M/300°, RFA1-MAND was swapped to SAI-02 due to IPA breaker trip-off, 4 minutes loss time. (Tinang MCR initiated)

22:00 - 23:00 SAI-01, 9720 kHz, 270M/285°, RFA1-CANT was swapped to SAI-02 due to IPA power supply troubleshooting.

Page 4 of 10 Attachment J. 7.docx

BROADCASTING BOARD OF GOVERNORS

INTERNATIONAL BROADCASTING BUREAU

Robert E. Kamosa Transmitting Station - Tinian

DAILY OPERATIONS LOG

Wednesday, January 21, 2015.

UTC TIME LOG ENTRY B14 BROADCAST SEASON

00:00:00 UTC START OF RADIO DAY

ON DUTY PERSONNEL CONTINUATION OF DAY SHIFT DUTY

Manny Sencio (MS), Edgard Acollador (EA), Edmund Degamo (edj), Ryan Diaz (RD)

BEGINNING OF RADIO DAY STATUS

ACTIVE DATABASE C:\schedule\B14-038.mdb

TRANSMITTERS All are operating properly.

ANTENNA & SLEW All are operating properly.

TOWER LIGHTS All are working properly.

REMARKS

MCR operators please upate Automation Intervention at the end of the log, if you have performed any manual operations.

Check for broadcast that needs to set frequency synthesizer to local mode.

Please note: When contacting anyone that needs to be informed/updated, it is advisable to fully identify yourself, and explain the nature of your call.

PENDING WORK

ORDERS

00:00:00 Next Scheduled Broadcast Change

INITIAL

TIN-02, on air @ 11695, 270L/270, NET RFA2 VIET.

Verified all frequencies and program output. RD

0023 Radioed RF to check if they are cleared from 267H. He informs they are cleared from the area and will be waiting for my call if it is locked out again.

0027 267H key returned to interlock panel.

0028 Reconfirmed 267H is cleared from RM.

00:30:00 Next Scheduled Broadcast Change

TIN-01, on air @ 15700, 267H/279, NET RFA2 BURM.

TIN-02, on air @ 17510, 280H/280, NET RFA2 BURM.

Verified all frequencies and program output. RD

01:00:00 Next Scheduled Broadcast Change

TIN-01, TIN-02, no changes.

0100-0200 TIN-07, subbed to 295H/295 @ 15105, 295M/295, NET RFA4 TIBE., due to 295M is locked out.

0100-0200 TIN-08, subbed to 333H/317 @ 17535, 313H/313, NET RFA3 UYGH., due to 313H is locked out.

Page 5 of 10 Attachment J. 7.docx

01:30:00 Next Scheduled Broadcast Change

INITIAL

TIN-01, TIN-02, off air.

TIN-07, TIN-08, no changes.

0132 267H antenna key secured. RF was informed.

02:00:00 Next Scheduled Broadcast Change

TIN-08, off air.

0200-0300 TIN-07, subbed to 295H/295 @ 15520, 295M/295, NET RFA4 TIBE., due to 295M is locked out.

03:00:00 Next Scheduled Broadcast Change

TIN-05, on air @ 21700, 333H/309, NET RFA1 MAND.

TIN-07, off air.

TIN-08, on air @ 11980, 295L/295, NET RFA1 MAND.

Verified all frequencies and program output. RD

0318 MS/edj schedules LOTO for 270L and 305L for verification of slew commands.

0320 MS/edj heads to phase 3 antenna field.

0329 TIN-05, off-on OVLD: Stages >N Defect indicated.

04:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 21560, 295H/303, NET VR1 CHIN.

TIN-05, no changes.

04:30:00 Next Scheduled Broadcast Change

0430 TIN-01, off air.

0450 RF returns 267H, 267L, and 280L.

0452 Emailed Tinang status of 267H, 267L, and 280L.

0459 RM returns 313H, 313L, and 295M.

0513 Emailed Tinang status of 313H, 313L, and 295M.

0520 MS/edj returned from phase 3 antenna field.

05:30:00 UTC - TRANSFER OVER OPERATIONS TO EVE SHIFT

ENDING STATUS Remains the same.

06:00:00 UTC - EVE SHIFT COMMENCE DUTY

Off Duty Personnel Manny Sencio (MS), Edgard Acollador (EA), Edmund Degamo (edj), Ryan Diaz

(RD)

On Duty Personnel Brian Hom (BH), Jeff Chant (JC), Iqbal Syed (IS)

BEGINNING STATUS C:\schedule\B14-038.mdb

Page 6 of 10 Attachment J. 7.docx

06:00:00 Next Scheduled Broadcast Change

INITIAL

TIN-01, on air @ 21480, 267H/279, NET RFA4 TIBE.

TIN-07, on air @ 15150, 280H/304, NET RFA1 MAND.

TIN-08, no changes.

Verified all frequencies and program output. JC

07:00:00 Next Scheduled Broadcast Change

TIN-01, TIN-05, TIN-07, off air.

Routine check of all transmitters.

0800 Phase 3 AC check OK @ 76 deg.

0930 TIN-03: Weekly, Quarterly SMP completed.

12:00:00 Next Scheduled Broadcast Change

TIN-02, on air @ 12055, 295M/295, NET RFA4 TIBE.

TIN-07, on air @ 05890, 333L/325 , NET T1 KORE.

12:30:00 Next Scheduled Broadcast Change

TIN-01, on air @ 11795, 280H/280, NET RFA2 BURM.

TIN-04, on air @ 13735, 313H/289, NET RFA2 BURM.

TIN-02, TIN-07, no change.

13:00:00 Next Scheduled Broadcast Change

TIN-02, on air @ 12050, 295M/295, NET RFA4 BURM.

TIN-08, on air @ 09360, 305L/329, NET RL-1 RU.

TIN-01,TIN-04,TIN-07, no change.

13:15:00 Next Scheduled Broadcast Change

TIN-05, on air @ 15470, 295H/271, NET VR2 VIET.

TIN-01,TIN-02,TIN-04,TIN-06,TIN-07, no change.

13:30:00 Next Scheduled Broadcast Change

TIN-04, on air @ 13735, 313H/289, NET RFA3 BURM.

TIN-02,TIN-05,TIN-07,TIN-08, no change.

TIN-01, off air.

14:00:00 UTC - REMOTE/UNMANNED OPERATION

Page 7 of 10 Attachment J. 7.docx

Off Duty Personnel Brian Hom (BH), Jeff Chant (JC), Iqbal Syed (IS)

On Duty Personnel Ambrose Ayuyu (AA)

14:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 12045, 280H/296, NET R CHIN.

TIN-02, on air @ 13655, 295M/287, NET RFA1 CANT.

TIN-05, on air @ 13870, 295H/271, NET RFA2 VIET.

TIN-08, @ 07235, 305L/329, NET T1 KORE.

TIN-07, no change.

TIN-04, off air.

1403 Called Tinang/Jordan techs off duty, leaving site.

15:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 06020, 280L/280, NET RFA1 MAND.

TIN-02, on air @ 09940, 295L/295, NET RFA4 TIBE.

TIN-04, on air @ 07210, 305L/329, NET RFA2 KORE.

TIN-07, on air @ 05855, 333L/333, NET RFA2 KORE.

TIN-08, on air @ 11985, 313L/321, NET RFA2 KORE.

TIN-05, off air.

1524 TIN-02, tripped off with UV DET. EXT. indication @ 9940, 295L/295, NET RFA4

TIBE. Immediately subbed to 270L.

1526-1600 TIN-02, subbed to 270L/294 @ 9940, 295L/295, NET RFA4 TIBE. 2 minutes lost time.

16:00:00 Next Scheduled Broadcast Change

TIN-02, on air @ 07415, 295L/303, NET RFA1 MAND.

TIN-04, TIN-07, TIN-08, no change.

16:30:00 Next Scheduled Broadcast Change

TIN-05, on air @ 09940, 270L/278, NET RFA4 BURM.

TIN-02, TIN-03,TIN-04,TIN-07, no change.

17:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 06020, 280L/296, NET RFA1 MAND.

TIN-02,on air @ 5415, 295L/303, TIN-05,TIN-07,no change.

TIN-04, TIN-08, off air.

17:30:00 Next Scheduled Broadcast Change

TIN-05, off air.

Page 8 of 10 Attachment J. 7.docx

TIN-01,TIN-02,TIN-07, no change.

18:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 6020, 280L/288, NET RFA1 MAND.

18:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 6020, 280L/288, NET RFA1 MAND.

19:00:00 Next Scheduled Broadcast Change

TIN-01, no change

TIN-02, off air.

TIN-07, on air @ 07415, 333L/317, NET RFA1 MAND.

TIN-08, on air @ 6095, 295L/303, NET RFA1 MAND.

20:00:00 Next Scheduled Broadcast Change

TIN-01, TIN-07, TIN-08, no changes.

TIN-05, on air @ 7495, 270L/278, NET RFA1 MAND.

21:00:00 Next Scheduled Broadcast Change

TIN-01, on air @ 11995, 333L/325, NET RFA2 KORE.

TIN-05, no changes.

TIN-07, on air @ 9690, 313L/329, NET RFA2 KORE.

TIN-08, on air @ 6020, 295L/303, NET RFA1 MAND.

2145 Ruben request for 305L/270L LOTO for 305L Vertical (sub) Slew Sw. P.M.

22:00:00 UTC - DAY SHIFT COMMENCE DUTY

Off Duty Personnel Ambrose Ayuyu (AA)

On Duty Personnel Manny Sencio (MS), Edgar Acollador (EA), Ryan Diaz (RD), Edmund Degamo Jr

(EDJ), Jess Domingo (JD)

22:00:00 Next Scheduled Broadcast Change

TIN-01, TIN-05, TIN-08, off air.

TIN-07, on air @ 15470, 333H//309, NET VR1 CHIN.

Verified all frequencies and program output. JD

2203 305L/270L locked out for B&G

2211 Called up TINANG/Nelson re. midshift status. He report TIN-02 IR at 1524UTC.

Took over control for Antenna LOTO

2213-2225 Prepare and activate “C:\Schedule\B14-038-Jan22.mdb:

2230 TIN-07, off air

Page 9 of 10 Attachment J. 7.docx

2245 Prepare and sent email to TINANG re. antenna LOTO for today

23:00:00 Next Scheduled Broadcast Change

TIN-01, off air.

TIN-07, on air @ 9825, 313L/305, NET RFA1 MAND.

2300-0000 TIN-08, sub to 333L/325 @ 11775, 305L/329, NET RFA1 MAND. 305L on LOTO

Verified all frequencies and program output. JD

23:15:00 Next Scheduled Broadcast Change

TIN-04, on air @ 15470, 295H/271, NET VR2 VIET.

TIN-07, TIN-08, no changes.

23:30:00 Next Scheduled Broadcast Change

2330-0000 TIN-02, sub to 280L/272 @ 11695, 270L/270, NET RFA2 VIET. 270L is on LOTO

TIN-04, TIN-07, TIN-08, no changes.

23:59:59 End of Radio Day

TRANSMITTER POWER, FILAMENT & PLATE METER READINGS

Transmitter Power Filament Plate OTHERS

TIN-01 10588620 27468.9 11717.7 SWITCHGEAR:

TIN-02 10530237 49459.02 60476.46 09162

TIN-03 04861646 20743.81 15474.15 CUC Meter:

TIN-04 02003378 37355.85 60499.57 7780

TIN-05 10393346 27339.40 11785.30 LOAD TAP:

TIN-06 4536074 27794.1 10532.50 279080

TIN-07 03058231 1504.8 6092.1 LT REPS:

TIN-08 2694711 12267.9 4610.1 28

FEEDER3 00225268 FEEDER4 00056973

DAILY TRANSMITTER BROADCAST AVAILABILITY

Transmitter Scheduled Broadcast Hours Lost Time (mins) Availability

TIN-01 10.500 0 100%

TIN-02 9.000 2 99.63%

TIN-03 0 0 100%

TIN-04 4.250 0 100%

TIN-05 8.750 0 100%

TIN-06 0 0 100%

TIN-07 14.500 0 100%

TIN-08 13.000 0 100%

DAILY TOTAL 56.000 2 99.944%

Page 10 of 10 Attachment J. 7.docx

DAILY OFF/ON REPORT

Transmitter Day Shift Eve Shift Mid Shift Daily Total

TIN-01

TIN-02

TIN-03

TIN-04

TIN-05

TIN-06

TIN-07

TIN-08

DAILY TOTAL 0 0 0 000

AUTOMATION INTERVENTION LOG

0023 Radioed RF to check if they are cleared from 267H. He informs they are cleared from the area and will be waiting for my call if it is locked out again.

0027 267H key returned to interlock panel.

0028 Reconfirmed 267H is cleared from RM.

0100-0200 TIN-07, subbed to 295H/295 @ 15105, 295M/295, NET RFA4 TIBE., due to 295M is locked out.

0100-0200 TIN-08, subbed to 333H/317 @ 17535, 313H/313, NET RFA3 UYGH., due to 313H is locked out.

0132 267H antenna key secured. RF was informed.

0200-0300 TIN-07, subbed to 295H/295 @ 15520, 295M/295, NET RFA4 TIBE., due to 295M is locked out.

0318 MS/edj schedules LOTO for 270L and 305L for verification of slew commands.

0320 MS/edj heads to phase 3 antenna field.

0329 TIN-05, off-on OVLD: Stages >N Defect indicated.

0450 RF returns 267H, 267L, and 280L.

0452 Emailed Tinang status of 267H, 267L, and 280L.

0459 RM returns 313H, 313L, and 295M.

0513 Emailed Tinang status of 313H, 313L, and 295M.

0520 MS/edj returned from phase 3 antenna field.

0800 Phase 3 AC check OK @ 76 deg.

0930 TIN-03: Weekly, Quarterly SMP completed.

1524 TIN-02, tripped off with UV DET. EXT. indication @ 9940, 295L/295, NET RFA4

TIBE. Immediately subbed to 270L.

1526-1600 TIN-02, subbed to 270L/294 @ 9940, 295L/295, NET RFA4 TIBE. 2 minutes lost time.

2203 305L/270L locked out for B&G

2211 Called up TINANG/Nelson re. midshift status. He report TIN-02 IR at 1524UTC.

Took over control for Antenna LOTO

2213-2225 Prepare and activate “C:\Schedule\B14-038-Jan22.mdb:

2230 TIN-07, off air

2245 Prepare and sent email to TINANG re. antenna LOTO for today

2300-0000 TIN-08, sub to 333L/325 @ 11775, 305L/329, NET RFA1 MAND. 305L on LOTO

2330-0000 TIN-02, sub to 280L/272 @ 11695, 270L/270, NET RFA2 VIET. 270L is on LOTO

File details come from the government source that posted it. Updated .